# Learning Distributions on Manifolds with Free-form Flows

### v1

@@ 7.3 Variance reduction
8 Experimental details
@@
In accordance with the details provided in Sec. 4.3, our approach incorporates multiple regularization loss components in addition to the negative log-likelihood objective. This results in the final loss expression:
8.1 Special orthogonal group
The special orthogonal group data set is synthetically generated. We refer to [10] for a description of the data generation process. They use an infinite stream of samples. To emulate this, we generate a data set of $10^{7}$ samples from their code, of which we reserve 1,000 for validation during training and 5,000 for testing. We vectorize the $3\times 3$ matrices before passing them into the fully-connected networks. All training details are given in Tab. 6.
8.2 Earth data
We follow previous works and use a dataset split of 80% for training, 10% for validation and 10% for testing. For the earth datasets we use a mixture of 5 learnable Von-Mises-Fisher distributions for the target latent distribution. We base our implementation on the hyperspherical_vae library [9]. In order to stabilize training we apply a small amount of Gaussian noise to every batch (see table 7) and project the resulting data point back onto the sphere. Other training hyperparameters can be found in table 8.
8.3 Tori
@@
The torus datasets are randomly split into a train dataset (80%), validation dataset (10%) and test dataset (10%). To counteract overfitting, we augment the RNA dataset with random Gaussian noise. The noise strength and total number of instances is reported in Tab. 9. We use a uniform latent distribution. While we train for $\sim$ 60k steps on the protein datasets (General, Glycine, Proline and Pre-Proline), we train for $\sim$ 120k steps on the RNA datasets, both with a batch size of 512. Further hyperparameters used in training can be found in Tab. 10.
8.4 Libraries
@@
We base our code on PyTorch [40], PyTorch Lightning [13], Lightning Trainable [25], Numpy [19], Matplotlib [21] for plotting and Pandas [31, 47] for data evaluation. We use the geomstats [32, 33] package for embeddings and projections.

### v2

@@ A.3 Variance reduction
Appendix B Experimental details
In accordance with the details provided in Section 4, our approach incorporates multiple regularization loss components in addition to the negative log-likelihood objective. This results in the final loss expression:
In all cases, we selected hyperparameter using the performance on the validation data.
B.1 Likelihood Evaluation
Sampling from a trained model can be easily achieved by sampling from the latent distribution and performing a single pass through the decoder $g$. However, in order to evaluate the likelihood of the test set under our model, as in (Draxler et al., 2024), we need to calculate the change of variable formula w.r.t. $g^{-1}$
$\log p_{X}(x)=\log p_{Z}(g^{-1}(x))+\log|R^{T}{g^{-1}}^{\prime}(x)Q|+\tfrac{1}{2}\log\frac{|R^{T}G_{Z}(g^{-1}(x))R|}{|Q^{T}G_{X}(x)Q|}$ (71)
$\approx\log p_{Z}(f(x))+\log|R^{T}g^{\prime}(f(x))^{-1}Q|+\tfrac{1}{2}\log\frac{|R^{T}G_{Z}(f(x))R|}{|Q^{T}G_{X}(x)Q|}.$ (72)
Here we used the approximation $f\approx g^{-1}$. While is expensive to compute for training, it is reasonably fast to compute during inference time. However, if the assumption of $f\approx g^{-1}$ is not sufficiently fulfilled the measured likelihoods might be inaccurate. We try to identify such cases by sampling points from $\mathcal{M}$ around the proposed latent point with small noise strength $\sigma$
$\tilde{z}=\pi(f(x)+\mathcal{N}(0,\sigma)).$ (73)
As inverse of the decoder $g$ we use the sample $\tilde{z}$ which results in the lowest reconstruction loss
$g^{-1}(x)\approx\underset{\tilde{z}}{\operatorname*{arg\,min}}||g(\tilde{z})-x||_{2}^{2}.$ (74)
We also do this with samples drawn around $\tilde{x}=\pi(x+\mathcal{N}(0,\sigma))$ and uniformly from the manifold $\tilde{z}=\mathcal{U}(\mathcal{M})$. We note that in most cases except for the earth datasets the likelihoods computed in this way agree with $f\approx g^{-1}$. In case of the earth datasets, we note that the newly computed likelihoods now agree with observed model quality. Specifically, whenever reconstruction loss is low we also see agreement between the likelihoods computed via $f$ and our approximation of $g^{-1}$ via sampling. Otherwise, the disagreement between $f$ and $g^{-1}$ and the resulting drop in model quality are correctly diagnosed. Therefore, for the all $\mathbb{S}^{2}$ experiments we report the likelihoods computed by our approximation.
B.2 Special orthogonal group
@@
To apply manifold free-form flows, we project an output matrix $R\in\mathbb{R}^{3\times 3}$ from the encoder/decoder to the subspace of special orthogonal matrices by finding the matrix $Q\in SO(3)$ minimizing the Frobenius norm $\|Q-R\|_{F}$. This is known as the constrained Procrustes problem and the solution $Q$ can be determined via the SVD $R=U\Sigma V^{T}$ (Lawrence et al., 2019):
$Q=USV^{T},$ (75)
The special orthogonal group data set is synthetically generated. We refer to (De Bortoli et al., 2022) for a description of the data generation process. They use an infinite stream of samples. To emulate this, we generate a data set of $10^{7}$ samples from their code, of which we reserve 1,000 for validation during training and 5,000 for testing. We vectorize the $3\times 3$ matrices before passing them into the fully-connected networks. All training details are given in Table 7, one training run takes approximately 7 hours on a NVIDIA A40.
The data set is synthetically generated; of the $N=100,000$ data points, we use 1% of for validation and hyperparameter selection and 5% for test NLL. Each run uses a different random initialization of papers.
B.3 Earth data
@@
We follow previous works and use a dataset split of 80% for training, 10% for validation and 10% for testing. For the earth datasets we use a mixture of 5 learnable Von-Mises-Fisher distributions for the target latent distribution. We base our implementation on the hyperspherical_vae library (Davidson et al., 2018). In order to stabilize training we apply a small amount of Gaussian noise to every batch (see table 8) and project the resulting data point back onto the sphere. Other training hyperparameters can be found in table 9. Each model trained around 20h on a compute cluster using a single NVIDIA A40.
B.4 Tori
@@
Figure 5: Log density of M-FFF models in the $(\Phi,\Psi)$-plane of protein backbone dihedral angles (known as Ramachandran plot(Ramachandran et al., 1963)). The learned density matches the true density indicated by the test dataset (black dots) very well. Note also that the learned distribution obeys the periodic boundary conditions.
@@
The torus datasets are randomly split into a train dataset (80%), validation dataset (10%) and test dataset (10%). To counteract overfitting, we augment the RNA dataset with random Gaussian noise. The noise strength and total number of instances is reported in Table 10. We use a uniform latent distribution. We train for  120k steps with a batch size of 512 which takes 2.5 to 3 hours on a NVIDIA GeForce RTX 2070 graphics card. Further hyperparameters used in training can be found in Table 11.
B.5 Hyperbolic space
A straightforward way to define distributions on hyperbolic space (but also other Riemannian manifolds) is, to define a probability density $p_{\text{tangent}}$ in the tangent space at the origin and use the exponential map $\exp_{0}$ to pushforward this distribution onto the manifold using Eq. 9:
$\log p_{\text{manifold}}(\exp_{0}(v))=\log p_{\text{tangent}}(v)-\log|J_{\exp_{0}}(v)|-\frac{1}{2}\log\frac{|G_{\text{manifold}}(\exp_{0}(v))|}{|G_{\text{tangent}}(v)|},$ (76)
where $G_{\text{manifold}}$ denotes the metric tensor of the embedded manifold and $G_{\text{tangent}}$ the metric tensor of the tangent space. This is also known as a ’wrapped’ distribution.
We use the Poincaré ball model, which embeds the n-dimensional hyperbolic space $\mathbb{H}^{n}$ in the n-dimensional Euclidean space $\mathbb{R}^{n}$ as defined in Table 2. The exponential map at the origin of this embedding and its Jacobian determinant are simply given by:
$\exp_{0}(v)=\tanh(\|v\|)\frac{v}{\|v\|}$ and $|J_{\exp_{0}}(v)|=\frac{\tanh(\|v\|)}{\|v\|\cosh^{2}(\|v\|)}.$ (77)
The metric tensor at some point $p\in\mathbb{H}$ is defined by: $G_{\mathbb{H}}^{ij}(p)=\lambda_{p}^{2}\delta_{ij}$ with $\lambda_{p}=2/(1-\|x\|^{2})$. The metric tensor of the tangent space is the usual Euclidean metric tensor: $G_{\mathbb{R}}^{ij}=\delta_{ij}$.
With this at hand, we can define latent and toy distributions as wrapped distributions at the origin, as depicted in Fig. 6.
For training, we sample 100k data points from each distribution. Hyperparameters for each model can be found in Table 12. Training takes approximately 2.5, 16, 8 and 16 hours on a NVIDIA A40 graphics card for the one Gaussian, five Gaussians, swish and checkerboard dataset respectively. The resulting model densities are shown in Fig. 6.
Figure 6: Density estimation on the Poincaré ball model. As latent distribution we use a wrapped normal distribution with standard deviation 0.5 (Left). As target distributions (top row) we define several toy distributions in the tangent space at the origin and use Eq. 76 to push forward to the manifold. We will reference each distribution from left to right as ’one Gaussian’, ’five Gaussians’, ’swish’ and ’checkerboard’. We train M-FFF on these target distributions using the full expression in Eq. 9 to compute the change in variables and evaluate the densities of the models (bottom row). M-FFF are capable to adapt to non-isometrically embedded manifolds. The learned densities on the one Gaussian, five Gaussians and swish dataset closely follow the target densities. On the checkerboard dataset, M-FFF cannot fully reproduce the sharp edges and density of the dataset.
Hyperparameter Value (one wrapped) Value (five gaussians) Value (swish) Value (checkerboard)
Layer type ResNet ResNet ResNet ResNet
residual blocks 2 6 6 6
Inner depth 2 3 3 3
Inner width $128$ $256$ $256$ $256$
Activation SiLU SiLU SiLU SiLU
$\beta_{\text{R}}^{x}$ $1000$ $1000$ $1000$ $1000$
$\beta_{\text{R}}^{z}$ $100$ $100$ $100$ $100$
$\beta_{\text{U}}^{x}$ $100$ $100$ $100$ $100$
$\beta_{\text{U}}^{z}$ 00 00 00 0
$\beta_{\text{P}}$ 1 1 1 1
Latent distribution Wrapped normal Wrapped normal Wrapped normal Wrapped Normal
Optimizer Adam Adam Adam Adam
Learning rate $2\times 10^{-4}$ $1\times 10^{-4}$ $1\times 10^{-4}$ $1\times 10^{-4}$
Scheduler Exponential w/ $\gamma=0.9986$ onecyclelr onecyclelr onecyclelr
Gradient clipping - - - -
Weight decay $1\times 10^{-3}$ $1\times 10^{-3}$ $1\times 10^{-3}$ $1\times 10^{-3}$
Batch size 4096 4096 4096 4096
Step count $\sim$ 84k $\sim$ 485k $\sim$ 240k $\sim$ 495k
Table 12: Details on the model architecture, loss weights and optimizer parameters for the Poincaré ball experiments.
B.6 3D mesh
Hyperparameter Value
Layer type ResNet
Residual blocks 2
Inner depth 5
Inner width $512$
Activation $\operatorname{ReLU}$
$\beta_{\text{R}}^{x}$ $1000$
$\beta_{\text{R}}^{z}$ 00
$\beta_{\text{U}}$ $10$
$\beta_{\text{P}}^{x}$ $100$
$\beta_{\text{P}}^{z}$ $10$
Latent distribution uniform
Optimizer Adam
Learning rate $5\times 10^{-4}$
Scheduler Exponential w/ $\gamma=1-0.0039$
Gradient clipping 1.0
Weight decay $3\times 10^{-5}$
Batch size 1,024
Step count 469.199
#Repetitions 3
Table 13: Hyperparameter choices for the bunny experiments.
We base our experiment on the manifold and data provided by (Chen and Lipman, 2024) using 80% for training, 10% for validation and hyperparameter tuning. We report test NLL on the remaining 10% of the data. Each run starts from different parameter initialization. They give the manifold as a triangular mesh, consisting of vertices $v_{i}\in\mathbb{R}^{3},i=1,\dots N_{v}$ and triangular faces $f_{j}\in\{1,\dots,N_{v}\}^{3}$.
Since the projection to the nearest point on the mesh has zero gradient in parts of $\mathbb{R}^{3}$, we instead project to the manifold using a separately trained auto encoder with a spherical latent space. This autoencoder consists of an encoder $e:\mathbb{R}^{3}\to\mathbb{R}^{3}$ consisting of five hidden layers with 256 neurons each, SiLU activations and an overall skip connection. The latent codes are computed by projecting the encoder outputs $e(x)$ to a sphere as $z(x)=e(x)/\|e(x)\|$, so that the latent space has the same the topology as the input mesh. Then, a decoder $d(z)$ with the same structure as the encoder is trained to reconstruct the original points by minimizing $\|x-d(e(x)/\|e(x)\|\|^{2}$. We train it for $1442^{18}=262,$ 144 steps, with each batch consisting of all $N_{v}=2,502$ vertices and an additional $N=2,502$ uniformly random points on the original mesh. We find that for successful training, it is helpful to filter out data with $x_{2}>0.5+n/10,000$, where $n$ is the step number. This prevents the long bunny ears from collapsing as the are slowly grown, allowing the model to adapt.
We then train M-FFF with the hyperparameters given in table Table 13, using the pretrained autoencoder as our projection to the manifold. Note that we do not train the distribution on the latent sphere of the encoder, but directly on the manifold spanned by it. Training takes approximately 14 hours on a NVIDIA A40.
B.7 Libraries
@@
We base our code on PyTorch (Paszke et al., 2019), PyTorch Lightning (Falcon and The PyTorch Lightning team, 2019), Lightning Trainable (Kühmichel and Draxler, 2023), Numpy (Harris et al., 2020), Matplotlib (Hunter, 2007) for plotting and Pandas (McKinney, 2010; The pandas development team, 2020) for data evaluation. We use the geomstats (Miolane et al., 2020, 2023) package for embeddings and projections.